OpenStudy (shal):

I need help with ONE math question!!!!1 Plzzz Help I will give medal!!!!!!! and be a fan!- Geometry

OpenStudy (shal):

12. Find the slope of the line that passes through (4, 1) and (7, –2). (Points : 5) –1 –1/3 1/3 1

OpenStudy (shal):

(7,-2)^^^ A.) -1 B. -1/3 ....

OpenStudy (anonymous):

\(\sf \Large m= \frac{y_2-y_1}{x_2-x_1}\), where m=slope , (4,1 ) =\((x_1, y_1)\) and (7,-2) = \((x_2, y_2)\) plug everything in the formula and you'll get it

OpenStudy (shal):

Can yu help me solve two more?

OpenStudy (anonymous):

did you get the answer for this question?

OpenStudy (shal):

im goin to solve it

OpenStudy (anonymous):

ok, solve that first before you move to the other one

OpenStudy (shal):

ok i said tht because i thought yu were goin to leave

OpenStudy (shal):

plz wait, thx for your time

OpenStudy (anonymous):

im just in the chat box , talking about naruto , so dont worry lol

OpenStudy (shal):

oh ok, thx, i got 1

OpenStudy (anonymous):

positive?

OpenStudy (shal):

yes

OpenStudy (anonymous):

i got negative ? O.o can you show me your solution?

OpenStudy (shal):

actually yes it is -1, i miscalculated it

OpenStudy (anonymous):

great :) so next one?

OpenStudy (shal):

let me attach it

OpenStudy (shal):

those two

OpenStudy (anonymous):

can you just take a screen shot? because i cannot download the file

OpenStudy (shal):

i did take screen shot

OpenStudy (anonymous):

save it as jpg not docx

OpenStudy (shal):

oh ok

OpenStudy (shal):

it wont work as jpg

OpenStudy (anonymous):

i can open it online so its okay. for the first one, do you know the formula for calculating distance between two points?

OpenStudy (shal):

no

OpenStudy (shal):

im so bad at math

OpenStudy (shal):

sorry, i like science better thts why

OpenStudy (anonymous):

oh nvm, you don't need it. i found a way how to do it. you have to use the formula for slope :) the one that we did before

OpenStudy (shal):

so tht formula?

OpenStudy (shal):

can yu help me solve it?

OpenStudy (anonymous):

don't worry, not everyone is good at everything, i suck at English/history etc. so for the first one.. calculate the slope of AB, BC, and AC using that formula

OpenStudy (shal):

can yu solve it fr me? because i dont feel like it im sick and i have cramps

OpenStudy (shal):

i mean the first part

OpenStudy (shal):

sorryyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyyy

OpenStudy (anonymous):

lol im very sleepy, i cant think properly we have to work this out together

OpenStudy (shal):

ok

OpenStudy (anonymous):

you'll solve for AB, i'll do BC

OpenStudy (anonymous):

Find the slope using the equation: (Y-Yo)=m(X-Xo) 1+2=m(4-7) 3=m(-3) m=-1 where m is the slope. Second question: ABC is NOT a right triangle because line segment AB != line segment CB where B is the point where they meet and count their displacement. A and C then say, "How far did you far?"... A Says, "3 units of dist." C says, "4 units of dist.". They have to be equal. Third Question: MidX = (x-xo)/2 = 1 MidY = (y-yo)/2=3 mid point (1,3) It's all about wondering how far you've gone.

OpenStudy (anonymous):

for the first one, YES IT IS A RIGHT TRIANGLE slope of AB = -3/4 slope of BC = 4/3 the negative reciprocal of one slope of a line means that it is perpendicular to that line.

OpenStudy (shal):

Thank you so much i just came back from vomiting.....

OpenStudy (anonymous):

vomiting because of math? O.o and thanks for sharing, i'm eating my ice cream

OpenStudy (shal):

no vomiting because im sick

OpenStudy (shal):

and i have so much homeowkr

OpenStudy (shal):

wait so 13 is d?

OpenStudy (anonymous):

that's okay :) you'll get over it. i have bunch of lab reports to do but i decided to relax for a bit. 13. NOPE. i just explained it

OpenStudy (shal):

and 14?

OpenStudy (anonymous):

13 is not D 14 here's how you would solve it: